Three times one number added to another number is 18. Twice the first number number minus the other number is 12. Find the numbers.
I ahve 5 and 9. Is this correct?

Answer by checkley75(3666)

3x+y=18 y=18-3x
2x-y=12 now substitute (18-3x) for y in this equation 7 solve for x
2x-(18-3x)=12
2x-18+3x=12
5x=30
x=30/5
x=6 answer now substitute 6 for x & solve for y
3*6+y=18
18+y=18
y=18-18
y=0 answer